Projection Admin 9/27

 To do this evening/ was completed: I'll get some content out of T3 tomorrow evening so we can free up some space for the NewFest features as we'll be teching them this Monday. 

Caught up with DMS for the T6/T2 inquiries. 

Made the rounds. We'll have to change the lamp in T3 sometime within the next week and a half ish. 

Presentation Issues:


Preshow Clips/Playlist: Wasn't sure DEBs got techd, so I took a look at it onscreen today before doors, adjusted volume. Looks/sounds great. 


Trailer Updates:


Ordering/Receiving (posters & trailers):


Ordering/Receiving (content): I unzipped and ingested Queerpanorama (Newfest feature) to the LMS today. 
I was trouble ingesting the DCP for Same, Again (also a NewFest feature). Downloading twice didn't work, so I checked it in DCP-o-matic since nothing seemed to be wrong with it at face value, and I only found the volume to be peaking, but again, sometimes a DCP will not ingest to the server if the volume is peaking too much. I adjusted the volume and it's now running again in DCP-o-matic, finishing to the preshow/trailers drive. Please ingest once done. 

The DCP of Paid in Full finished in T1. It's copying over to the George Ivan drive. Please ingest once done. 


Outgoing/Shipping (content): Handed off Cyborg and Trolls 2 to FedEx this morning. 


Print/video Inspection: 


Asset tracking doc updated/checked?: 


Maintenance/cleaning: 


Cue sheets updated/checked (every shift): Yes

Marquee updated?: 

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:







Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will